# CLIPin: A Non-contrastive Plug-in to CLIP for Multimodal Semantic Alignment

### v1

@@ Introduction
CLIP has shown remarkable success in learning joint representations from large-scale image-text pairs, delivering strong performance across a wide range of downstream tasks in both natural and medical domains (Radford et al. 2021; Jia et al. 2021; Goel et al. 2022; Zhang et al. 2022b; Huang et al. 2021; Du et al. 2024). Despite its effectiveness, CLIP often suffers from inherent challenges in image-text datasets. Specifically, many large-scale natural image-text datasets used in CLIP-style pretraining (Thomee et al. 2016; Sharma et al. 2018; Schuhmann et al. 2021) are automatically crawled from the web with minimal or no human supervision, resulting in loose or inaccurate aligned pairs. This semantic noise undermines effective cross-modal representation learning by introducing ambiguity, where a single image or caption may be partially relevant to multiple samples within a batch (Zhou et al. 2023; Li et al. 2021a, 2022; Jia et al. 2021; Wu et al. 2022). For medical datasets, they usually exhibit accurate alignment, since the reports are written by clinicians based on image readings. However, the diversity of textual descriptions is limited due to the small variety of diseases and anatomical variations. In these cases, the CLIP often suffers from semantically similar samples being treated as negative sample pairs (negatives) (Yang et al. 2024; Wang et al. 2022). Although these two issues differ in form (semantic looseness in natural datasets and semantic redundancy in medical datasets), they both violate the core assumption of the InfoNCE loss (Oord, Li, and Vinyals 2018), namely that each positive pair is surrounded by mutually exclusive negatives. As a result, the model supervision becomes noisy or ambiguous, ultimately impairing the quality of learned representations.
@@
Prior works have attempted to enhance representation quality under these limitations by introducing architectural modifications and multi-task objectives, such as incorporating image-text matching (ITM) losses and cross-modal attention mechanisms (Li et al. 2021a, 2022). While these methods introduce complex constraints, they are grounded in the contrastive learning paradigm, thus inherit its limitations. Other approaches have incorporated non-contrastive components to improve inter-modal alignment and intra-modal diversity from a distributional perspective (Zhou et al. 2023). However, they typically lack explicit modeling of fine-grained, instance-level semantic correspondence.
@@
To address these challenges, we propose CLIPin, a unified plug-in that enables non-contrastive feature representation to integrate with CLIP-style architectures, to enhance multimodal representation learning within image-text pretraining paradigms. Our key contributions are as follows: (i) We introduce a general and modular non-contrastive strategy that can be seamlessly integrated into existing contrastive frameworks without modifying their base architectures. By leveraging two semantically consistent yet independently augmented views per sample, our approach enables diverse and robust representation learning through distinct pathways without additional supervision. (ii) We design two shared pre-projectors for image and text modalities respectively, for facilitating the integration of contrastive and non-contrastive branches in a parameter-compromise manner. (iii) Extensive experiments across a wide range of downstream tasks demonstrate that CLIPin consistently improves feature quality and cross-modal alignment, while serving as a plug-and-play module with strong generalizability across various contrastive architectures.
Related work
@@
Contrastive learning was first established in single-modal representation learning, particularly in vision tasks. Methods such as (Caron et al. 2021; Oquab et al. 2024; Chen et al. 2020; Caron et al. 2020; Li et al. 2021b) have achieved impressive performance by contrasting different augmented views of the same image and learning inter-instance discrimination. Despite its simplicity and effectiveness, contrastive learning still faces practical challenges, particularly its heavy reliance on both the quantity and quality of negative sample pairs. On the one hand, effective estimation of the InfoNCE objective requires large batch sizes, which imposes significant memory and hardware demands. On the other hand, the representativeness and semantic diversity of negative sample pairs are crucial, unrepresentative or semantically similar negatives can reduce alignment precision and impair training. To address these limitations, methods like MoCo (He et al. 2020) introduce a memory bank and momentum encoder to decouple batch size from the number of negatives. Other approaches, such as PCL (Li et al. 2021b) and SwAV (Caron et al. 2020), employ clustering to avoid semantically redundant negatives, thereby improving training stability and representation quality.
Building on the success of vision-only models, contrastive learning has become a dominant paradigm in multimodal representation learning, with CLIP (Radford et al. 2021) as a representative framework. CLIP adopts a dual-encoder architecture trained with InfoNCE loss to align image and text representations in a shared embedding space. By pulling features of paired samples together and pushing mismatched ones apart, CLIP enables significant performance across diverse downstream tasks in both natural and medical domains. To improve robustness in the multimodal setting, recent works have augmented contrastive frameworks with auxiliary objectives (e.g., image-text matching, masked language modeling, or caption generation) and architectural refinements such as momentum encoders and query-based transformers (Li et al. 2021a, 2022, 2023; Yu et al. 2022).
@@
Non-contrastive learning offers a compelling alternative by eliminating the need for negative sample pairs (Grill et al. 2020; Chen and He 2021; Zbontar et al. 2021; Jing et al. 2022; Wen and Li 2022). Methods such as SimSiam (Chen and He 2021) and BYOL (Grill et al. 2020) achieve representation learning by encouraging consistency between positive pairs (e.g., different augmentations of the same sample) using an online-target architecture, where the target network is updated via exponential moving average (EMA). These approaches have shown strong performance in single-modal tasks, but their adoption in multimodal settings remains limited, because non-contrastive methods are highly sensitive to the interplay between model capacity and data scale, relying heavily on strong augmentations, and requiring careful design to avoid representation collapse (Li, Efros, and Pathak 2022; Wetzer, Lindblad, and Sladoje 2023; Vahidi et al. 2024; Huang et al. 2024; Wen and Li 2022; Zhang et al. 2022a). In multimodal contexts, where image and text encoders are inherently heterogeneous, these issues are further amplified.
Until now, only xCLIP (Zhou et al. 2023) has attempted to extend non-contrastive learning to vision-language settings, which aligns the output distributions of the image and text encoders by optimizing both their sharpness and smoothness. However, its non-contrastive component focuses solely on batch-level distribution alignment and lacks explicit modeling of instance-level semantic correspondence. Furthermore, its training objective is decoupled from CLIP-style representation learning, limiting its compatibility with existing contrastive frameworks and weakening the interpretability of learned alignments.
Method
Figure 1: Overview of the proposed CLIPin framework. (a) Typical contrastive text-image pretraining architecture (i.e., CLIP). (b) CLIPin architecture with key modules, loss functions, and parameter update strategy. (c) Detailed structure of projectors and predictors in CLIPin.
Non-Contrastive multimodal structure of CLIPin
@@
To address the limitations of CLIP in learning robust and generalizable representations, particularly its vulnerability to semantic looseness and redundancy, we propose CLIPin, a unified non-contrastive plug-in that can be seamlessly integrated into CLIP-style architectures to enhance cross-modal semantic alignment, inspired by momentum-based dual-branch architectures in self-supervised learning (Grill et al. 2020). Unlike the original CLIP framework illustrated in Fig. 1(a), which relies exclusively on contrastive learning with negative sample pairs, CLIPin incorporates a non-contrastive pathway built on a symmetric online–target architecture for both image and text modalities. This results in parallel processing branches that facilitate both inter- and intra-modal alignment jointly. Each branch includes a modality-specific encoder, a projector, and a predictor (only on the online side). The target branch omits the predictor to introduce asymmetry and is updated via exponential moving average (EMA) of the corresponding online branch.
@@
We now describe the architecture of CLIPin in detail, as illustrated in Fig. 1(b). For each image-text pair in a training batch, the input image is augmented by two random transformations of equal strength, producing $I^{(1)}$ and $I^{(2)}\in\mathbb{R}^{3\times H\times W}$. The corresponding text $T$ is tokenized and augmented to obtain $\hat{T}^{(1)}$ and $\hat{T}^{(2)}\in\mathbb{R}^{l}$, where $l$ denotes the maximum text length.
@@
$\mathcal{M}_{\text{online},\mathrm{I/T}}(\cdot)$ $=g_{\mathrm{I/T}}\big{(}f_{\theta/\phi}(\cdot)\big{)},$ (1)
$\mathcal{M}_{\text{target},\mathrm{I/T}}(\cdot)$ $=g_{\mathrm{I/T}}^{\mathrm{m}}\big{(}f_{\theta/\phi}^{\mathrm{m}}(\cdot)\big{)},$
where $f_{\theta/\phi}$ denotes the image or text encoder, and $g_{\mathrm{I/T}}$ is the corresponding modality-specific projector, which will be elaborated in Section Contrastive learning from shared pre-projectors. The momentum versions, $f_{\theta/\phi}^{\mathrm{m}}$ and $g_{\mathrm{I/T}}^{\mathrm{m}}$, constitute the target branches. Parameters in the target branches are updated using an EMA of the online parameters:
$\quad\;\;\mathcal{M}_{\text{target},\mathrm{I/T}}^{0}=\mathcal{M}_{\text{online},\mathrm{I/T}}^{0},$ (2)
@@
where $t$ is the training step and $\beta$ is the momentum coefficient. $q_{\mathrm{inter,I}}$ and $q_{\mathrm{inter,T}}$ are the image and text predictors that appended to the online branches to introduce asymmetry that helps prevent collapse (Grill et al. 2020; Chen and He 2021). The predicted features from the online branches are:
$\mathbf{u}$ $=q_{\mathrm{inter,I}}\big{(}\mathcal{M}_{\text{online},\mathrm{I}}(I^{(1)})\big{)},$ (3)
$\mathbf{v}$ $=q_{\mathrm{inter,T}}\big{(}\mathcal{M}_{\text{online},\mathrm{T}}(\hat{T}^{(1)})\big{)}.$
@@
$\mathbf{u}_{\mathrm{intra}}$ $=q_{\mathrm{intra,I}}\big{(}\mathcal{M}_{\text{online},\mathrm{I}}(I^{(1)})\big{)},$ (6)
$\mathbf{v}_{\mathrm{intra}}$ $=q_{\mathrm{intra,T}}\big{(}\mathcal{M}_{\text{online},\mathrm{T}}(\hat{T}^{(1)})\big{)}.$
@@
Contrastive learning from shared pre-projectors
@@
Although CLIPin is a non-contrastive plug-in specifically designed to be integrated with contrastive learning in a single framework, its architectural requirements, especially the projectors, differ from those of conventional contrastive learning. While it is conceivable that a shared projector could support both paradigms, practical considerations often call for distinct designs. Empirical evidence (Chen and He 2021; Zhou et al. 2023) suggests that non-contrastive methods typically rely on more complex projector designs, characterized by deeper architectures and higher output dimensionalities. In contrast, contrastive methods favor simpler and lower-dimensional projectors. For example, CLIP reduces encoder output to 512 dimensions via a linear layer, whereas non-contrastive approaches like SimSiam project features to 2,048 dimensions using a multi-layer perceptron (MLP). More notably, xCLIP (Zhou et al. 2023) expands the encoder output to 32,768 dimensions through a bottleneck module to achieve optimal performance.
This divergence arises from the different roles of projectors in each paradigm. In contrastive learning, the projector acts as an ”information bottleneck”, preserving only essential semantic content while discarding irrelevant details. This supports the alignment of semantically related image-text pairs and the separation of unrelated ones. A high-dimensional projector may capture excessive nuisance signals, hindering generalization across modalities (Gupta et al. 2022; Ouyang et al. 2025; Huang et al. 2024; Jing et al. 2022). In contrast, non-contrastive learning does not rely on negative sample pairs, making it less sensitive to overfitting noise in high-dimensional spaces. In this case, higher-dimensional representations can be beneficial for capturing fine-grained features and improving the overall performance. Moreover, deeper projector networks help mitigate representation collapse, a known limitation of non-contrastive objectives.
@@
To integrate contrastive and non-contrastive learning for enhanced representation quality, we design the projectors $(g_{\mathrm{I/T}},g_{\mathrm{I/T}}^{\mathrm{m}})$ and predictors $(q_{\mathrm{intra,I/T}},q_{\mathrm{inter,I/T}})$ as bottleneck, drawing inspiration from (Zhou et al. 2023; Chen and He 2021), and decompose each projector into two components: (i) a shared pre-projector $(g_{\mathrm{pre,I/T}},g_{\mathrm{pre,I/T}}^{\mathrm{m}})$, and (ii) a CLIPin-specific sub-projector $(g_{\mathrm{NCL,I/T}},g_{\mathrm{NCL,I/T}}^{\mathrm{m}})$, as illustrated in Fig. 1(c). After this decomposition, the online and target branches for the image and text modalities are structured as:
@@
$\mathbf{u}_{\mathrm{CL}}$ $=g_{\mathrm{CL,I}}\big{(}g_{\mathrm{pre,I}}\big{(}f_{\theta}(I^{(1)})\big{)}\big{)},$ (9)
$\mathbf{v}_{\mathrm{CL}}$ $=g_{\mathrm{CL,T}}\big{(}g_{\mathrm{pre,T}}\big{(}f_{\phi}(\hat{T}^{(1)})\big{)}\big{)},$
@@
$\mathbf{U}_{\mathrm{CL}}=\{\mathbf{u}_{\mathrm{CL},1},\dots,\mathbf{u}_{\mathrm{CL},B}\},\,\mathbf{V}_{\mathrm{CL}}=\{\mathbf{v}_{\mathrm{CL},1},\dots,\mathbf{v}_{\mathrm{CL},B}\},$ (10)
@@
$\mathcal{L}_{\mathrm{CL,I2T}}=$ (11)
$-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\big{(}\operatorname{Norm}(\mathbf{u}_{\mathrm{CL},i})^{\top}\operatorname{Norm}(\mathbf{v}_{\mathrm{CL},i})/\tau\big{)}}{\sum_{j=1}^{B}\exp\big{(}\operatorname{Norm}(\mathbf{u}_{\mathrm{CL},i})^{\top}\operatorname{Norm}(\mathbf{v}_{\mathrm{CL},j})/\tau\big{)}},$
$\mathcal{L}_{\mathrm{CL,T2I}}=$
$-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\big{(}\operatorname{Norm}(\mathbf{v}_{\mathrm{CL},i})^{\top}\operatorname{Norm}(\mathbf{u}_{\mathrm{CL},i})/\tau\big{)}}{\sum_{j=1}^{B}\exp\big{(}\operatorname{Norm}(\mathbf{v}_{\mathrm{CL},i})^{\top}\operatorname{Norm}(\mathbf{u}_{\mathrm{CL},j})/\tau\big{)}},$
$\mathcal{L}_{\mathrm{CL}}=\mathcal{L}_{\mathrm{CL,I2T}}+\mathcal{L}_{\mathrm{CL,T2I}}.$
@@
Experiments
Experiment settings

### v2

@@ 1 Introduction
CLIP has shown remarkable success in learning joint representations from large-scale image-text pairs, delivering strong performance across a wide range of downstream tasks in both natural and medical domains [1, 2, 3, 4, 5, 6]. Despite its effectiveness, CLIP often suffers from inherent challenges in image-text datasets. Specifically, many large-scale natural image-text datasets used in CLIP-style pretraining [7, 8, 9] are automatically crawled from the web with minimal or no human supervision, resulting in loose or inaccurate aligned pairs. This semantic noise undermines effective cross-modal representation learning by introducing ambiguity, where a single image or caption may be partially relevant to multiple samples within a batch [10, 11, 12, 2, 13]. For medical datasets, they usually exhibit accurate alignment, since the reports are written by clinicians based on image readings. However, the diversity of textual descriptions is limited due to the small variety of diseases and anatomical variations. In these cases, the CLIP often suffers from semantically similar samples being treated as negative sample pairs (negatives) [14, 15]. Although these two issues differ in form (semantic looseness in natural datasets and semantic redundancy in medical datasets), they both violate the core assumption of the InfoNCE loss [16], namely that each positive pair is surrounded by mutually exclusive negatives. As a result, the model supervision becomes noisy or ambiguous, ultimately impairing the quality of learned representations.
@@
Prior works have attempted to enhance representation quality under these limitations by introducing architectural modifications and multi-task objectives, such as incorporating image-text matching (ITM) losses and cross-modal attention mechanisms [11, 12]. While these methods introduce complex and effective constraints, they remain grounded in the contrastive learning paradigm, thus inherit its limitations. Other approaches have incorporated non-contrastive components to improve inter-modal alignment and intra-modal diversity from a distributional perspective [10]. However, they typically lack explicit modeling of fine-grained, instance-level semantic correspondence.
@@
To address these challenges, we propose CLIPin, a unified plug-in that enables non-contrastive feature representation to integrate with CLIP-style architectures, to enhance multimodal representation learning within image-text pretraining paradigms. Our key contributions are as follows: (i) We introduce a general and modular non-contrastive strategy that can be seamlessly integrated into existing contrastive frameworks without modifying their base architectures. By leveraging two semantically consistent yet independently augmented views per sample, our approach enables diverse and robust representation learning through distinct pathways without additional supervision. (ii) We design two shared pre-projectors for image and text modalities respectively, for facilitating the integration of contrastive and non-contrastive branches in a parameter-compromise manner. (iii) Extensive experiments across a wide range of downstream tasks, demonstrate that CLIPin consistently improves feature quality and cross-modal alignment, while serving as a plug-and-play module with strong generalizability across various contrastive architectures.
2 Related work
@@
Contrastive learning was first established in single-modal representation learning, particularly in vision tasks. Methods such as [17, 18, 19, 20, 21] have achieved impressive performance by contrasting different augmented views of the same image and learning inter-instance discrimination. Despite its simplicity and effectiveness, contrastive learning still faces practical challenges, particularly its heavy reliance on both the quantity and quality of negative sample pairs. On the one hand, effective estimation of the InfoNCE objective requires large batch sizes, which imposes significant memory and hardware demands. On the other hand, the representativeness and semantic diversity of negative sample pairs are crucial, unrepresentative or semantically similar negatives can reduce alignment precision and impair training. To address these limitations, methods like MoCo [22] introduce a memory bank and momentum encoder to decouple batch size from the number of negatives. Other approaches, such as PCL [21] and SwAV [20], employ clustering to avoid semantically redundant negatives, thereby improving training stability and representation quality.
Building on the success of vision-only models, contrastive learning has become a dominant paradigm in multimodal representation learning, with CLIP [1] as a representative framework. CLIP adopts a dual-encoder architecture trained with InfoNCE loss to align image and text representations in a shared embedding space. By pulling features of paired samples together and pushing mismatched ones apart, CLIP enables significant performance across diverse downstream tasks in both natural and medical domains. To improve robustness in the multimodal setting, recent works have augmented contrastive frameworks with auxiliary objectives (e.g., image-text matching, masked language modeling, or caption generation) and architectural refinements such as momentum encoders and query-based transformers [11, 12, 23, 24].
@@
Non-contrastive learning offers a compelling alternative by eliminating the need for negative sample pairs [25, 26, 27, 28, 29]. Methods such as SimSiam [26] and BYOL [25] achieve representation learning by encouraging consistency between positive pairs (e.g., different augmentations of the same sample) using an online-target architecture, where the target network is updated via exponential moving average (EMA). These approaches have shown strong performance in single-modal tasks, but their adoption in multimodal settings remains limited, because non-contrastive methods are highly sensitive to the interplay between model capacity and data scale, relying heavily on strong augmentations, and requiring careful design to avoid representation collapse [30, 31, 32, 33, 29, 34]. In multimodal contexts, where image and text encoders are inherently heterogeneous, these issues are further amplified.
Until now, only xCLIP [10] has attempted to extend non-contrastive learning to vision-language settings, which aligns the output distributions of the image and text encoders by optimizing both their sharpness and smoothness. However, its non-contrastive component focuses solely on batch-level distribution alignment and lacks explicit modeling of instance-level semantic correspondence. Furthermore, its training objective is decoupled from CLIP-style representation learning, limiting its compatibility with existing contrastive frameworks and weakening the interpretability of learned alignments.
3 Method
Figure 1: Overview of the proposed CLIPin framework. (a) CLIPin architecture with key modules, loss functions, and parameter update strategy. (b) Detailed structure of projectors and predictors in CLIPin.
3.1 Non-Contrastive multimodal architecture of CLIPin
@@
To address the limitations of CLIP in learning robust and generalizable representations, particularly its vulnerability to semantic looseness and redundancy, we propose CLIPin, a unified non-contrastive plug-in that can be seamlessly integrated into CLIP-style architectures to enhance cross-modal semantic alignment, inspired by momentum-based dual-branch architectures in self-supervised learning [25]. Unlike the original CLIP framework, which relies exclusively on contrastive learning with negative sample pairs, CLIPin incorporates a non-contrastive pathway built on a symmetric online–target architecture for both image and text modalities. This results in parallel processing branches that facilitate both inter- and intra-modal alignment jointly. Each branch includes a modality-specific encoder, a projector, and a predictor (only on the online side). The target branch omits the predictor to introduce asymmetry and is updated via exponential moving average (EMA) of the corresponding online branch.
@@
We now describe the architecture of CLIPin in detail, as illustrated in Fig. 1(a). For each image-text pair in a training batch, the input image is augmented by two random transformations of equal strength, producing $I^{(1)}$ and $I^{(2)}\in\mathbb{R}^{3\times H\times W}$. The corresponding text $T$ is tokenized and augmented to obtain $\hat{T}^{(1)}$ and $\hat{T}^{(2)}\in\mathbb{R}^{l}$, where $l$ denotes the maximum text length.
@@
$\mathcal{M}_{\text{online},\mathrm{I/T}}(\cdot)=g_{\mathrm{I/T}}\big(f_{\theta/\phi}(\cdot)\big),\quad\mathcal{M}_{\text{target},\mathrm{I/T}}(\cdot)=g_{\mathrm{I/T}}^{\mathrm{m}}\big(f_{\theta/\phi}^{\mathrm{m}}(\cdot)\big),$ (1)
where $f_{\theta/\phi}$ denotes the image or text encoder, and $g_{\mathrm{I/T}}$ is the corresponding modality-specific projector, which will be elaborated in Section 3.2. The momentum versions, $f_{\theta/\phi}^{\mathrm{m}}$ and $g_{\mathrm{I/T}}^{\mathrm{m}}$, constitute the target branches. Parameters in the target branches are updated using an EMA of the online parameters:
$\mathcal{M}_{\text{target},\mathrm{I/T}}^{0}$ $=\mathcal{M}_{\text{online},\mathrm{I/T}}^{0},$ (2)
@@
where $t$ is the training step and $\beta$ is the momentum coefficient. $q_{\mathrm{inter,I}}$ and $q_{\mathrm{inter,T}}$ are the image and text predictors that appended to the online branches to introduce asymmetry that helps prevent collapse [25, 26]. The predicted features from the online branches are:
$\mathbf{u}=q_{\mathrm{inter,I}}\big(\mathcal{M}_{\text{online},\mathrm{I}}(I^{(1)})\big),\quad\mathbf{v}=q_{\mathrm{inter,T}}\big(\mathcal{M}_{\text{online},\mathrm{T}}(\hat{T}^{(1)})\big).$ (3)
@@
$\mathbf{u}_{\mathrm{intra}}=q_{\mathrm{intra,I}}\big(\mathcal{M}_{\text{online},\mathrm{I}}(I^{(1)})\big),\quad\mathbf{v}_{\mathrm{intra}}=q_{\mathrm{intra,T}}\big(\mathcal{M}_{\text{online},\mathrm{T}}(\hat{T}^{(1)})\big).$ (6)
@@
3.2 Contrastive learning from shared pre-projectors
@@
Although CLIPin is a non-contrastive plug-in specifically designed to be integrated with contrastive learning in a single framework, its architectural requirements, especially the projectors, differ from those of conventional contrastive learning. While it is conceivable that a shared projector could support both paradigms, practical considerations often call for distinct designs. Empirical evidence [26, 10] suggests that non-contrastive methods typically rely on more complex projector designs, characterized by deeper architectures and higher output dimensionalities. In contrast, contrastive methods favor simpler and lower-dimensional projectors. For example, CLIP reduces encoder output to 512 dimensions via a linear layer, whereas non-contrastive approaches like SimSiam project features to 2,048 dimensions using a multi-layer perceptron (MLP). More notably, xCLIP [10] expands the encoder output to 32,768 dimensions through a bottleneck module to achieve optimal performance.
This divergence arises from the different roles of projectors in each paradigm. In contrastive learning, the projector acts as an "information bottleneck", preserving only essential semantic content while discarding irrelevant details. This supports the alignment of semantically related image-text pairs and the separation of unrelated ones. A high-dimensional projector may capture excessive nuisance signals, hindering generalization across modalities [35, 36, 33, 28]. In contrast, non-contrastive learning does not rely on negative sample pairs, making it less sensitive to overfitting noise in high-dimensional spaces. In this case, higher-dimensional representations can be beneficial for capturing fine-grained features and improving the overall performance. Moreover, deeper projector networks help mitigate representation collapse, a known limitation of non-contrastive objectives.
@@
To integrate contrastive and non-contrastive learning for enhanced representation quality, we design the projectors $(g_{\mathrm{I/T}},g_{\mathrm{I/T}}^{\mathrm{m}})$ and predictors $(q_{\mathrm{intra,I/T}},q_{\mathrm{inter,I/T}})$ as bottleneck, drawing inspiration from [10, 26], and decompose each projector into two components: (i) a shared pre-projector $(g_{\mathrm{pre,I/T}},g_{\mathrm{pre,I/T}}^{\mathrm{m}})$, and (ii) a CLIPin-specific sub-projector $(g_{\mathrm{NCL,I/T}},g_{\mathrm{NCL,I/T}}^{\mathrm{m}})$, as illustrated in Fig. 1(b). After this decomposition, the online and target branches for the image and text modalities are structured as:
@@
$\mathbf{u}_{\mathrm{CL}}=g_{\mathrm{CL,I}}\big(g_{\mathrm{pre,I}}\big(f_{\theta}(I^{(1)})\big)\big),\quad\mathbf{v}_{\mathrm{CL}}=g_{\mathrm{CL,T}}\big(g_{\mathrm{pre,T}}\big(f_{\phi}(\hat{T}^{(1)})\big)\big),$ (9)
@@
$\mathbf{U}_{\mathrm{CL}}=\{\mathbf{u}_{\mathrm{CL},1},\dots,\mathbf{u}_{\mathrm{CL},B}\},\quad\mathbf{V}_{\mathrm{CL}}=\{\mathbf{v}_{\mathrm{CL},1},\dots,\mathbf{v}_{\mathrm{CL},B}\},$ (10)
@@
$\mathcal{L}_{\mathrm{CL,I2T}}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\big(\operatorname{Norm}(\mathbf{u}_{\mathrm{CL},i})^{\top}\operatorname{Norm}(\mathbf{v}_{\mathrm{CL},i})/\tau\big)}{\sum_{j=1}^{B}\exp\big(\operatorname{Norm}(\mathbf{u}_{\mathrm{CL},i})^{\top}\operatorname{Norm}(\mathbf{v}_{\mathrm{CL},j})/\tau\big)},$ (11)
$\mathcal{L}_{\mathrm{CL,T2I}}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\big(\operatorname{Norm}(\mathbf{v}_{\mathrm{CL},i})^{\top}\operatorname{Norm}(\mathbf{u}_{\mathrm{CL},i})/\tau\big)}{\sum_{j=1}^{B}\exp\big(\operatorname{Norm}(\mathbf{v}_{\mathrm{CL},i})^{\top}\operatorname{Norm}(\mathbf{u}_{\mathrm{CL},j})/\tau\big)},$
$\quad\quad\quad\quad\quad\mathcal{L}_{\mathrm{CL}}=\mathcal{L}_{\mathrm{CL,I2T}}+\mathcal{L}_{\mathrm{CL,T2I}}.$
@@
4 Experiments
4.1 Experiment settings